# Sequential Conformal Risk Control UQ Uncertainty Quantification

### v1

@@ B-B A simple symmetry property
Appendix C Details on Experiments
@@
C-A Complete benchmark
We include the complete results of our experiments here in Tables V, VI, VII, and VIII. We now introduce multiple additional visualizations of our best-performing experiment run, for additional discussions on the limitations of conformal object detection. On Fig. 9, we observe two successful detections, with singleton classification prediction sets and small localization margins. However, on the third image, the moon is predicted to be a kite. While this is incorrect, due to the monotonicity constraints of CRC, we may only provide recall guarantees, i.e., ensure that we do not miss the ground truth. In that framework, additional predictions such as the moon as a kite here, are not penalized. In Fig. 10, we observe another instance of the same mistake, however, here the boat is detected twice, once excluding, and once including its mast. We however, observe two failure cases of our guarantee, corresponding to previously discussed issues. Indeed, in both examples, there are abuses of labeling: on one books are sometimes annotated individually, and sometimes as a block. On the second, a statue is annotated as a person. The variability in the definition of ground truths leads to perceived failures, that are not necessarily actual failures. As visible on both images, the clock and statue (person) are respectively assigned a book and a clock predictions by our matching. In order to reach lower levels of error rates $\alpha$, these mistakes would have to be controlled, and would result in a need for significant margins at inference (or models that are able to learn this bias).
C-B Additional Experiments
@@
C-B1 Monotonicty
In this section, we conduct an additional study on the monotonicity of loss functions, used or considered at a point in our work.
C-B2 mAP vs CP
@@ C-B3 Margins
N. margins Value of the margins Cov Set Size
@@
Appendix D Effect of Object Detection and Conformal Settings
@@
Lastly, regarding the $x_{1}y_{1}x_{2}y_{2}$ (coordinates of two corners) versus $xywh$ (coordinates of the center + width and height) bounding box representations, while fundamentally equivalent, their application within CP algorithms yields divergent results. The $xywh$ format inherently involves symmetric corrections and a translational adjustment, whereas the $xyxy$ format necessitates independent corrections for each side. However, these approaches can be reformulated to yield equivalent predictive sets.
Appendix E Additional Losses
As discussed in details in Appendix A, many natural image-level losses are aggregations of object-level losses. While there are exceptions to that, the losses presented in this appendix are of this form. Formally, denote by $\Phi:S^{\infty}(\mathbb{R})\rightarrow\mathbb{R}$ an aggregation function, mapping a sequence $S$ of arbitrary length – with $S_{i}$ valued in $\mathbb{R}$ – to $\mathbb{R}$. We introduce three main aggregation strategies.
  1. 1.
  2. 2.
    Max: $\Phi_{\mathrm{max}}(S)=\max_{i}S_{i}$
  3. 3.
    Thresholded Average (at level $\tau$): $\Phi_{\mathrm{thr}}(S)=\mathbbm{1}_{\Phi_{\mathrm{thr}}(S)>\tau}$
E-A Confidence
where $d:(\mathcal{B}\times\mathcal{C})\times(\mathcal{B}\times\Sigma_{K-1})\rightarrow\mathbb{R}^{+}$ is any function measuring a distance (particularly, it does not need to be symmetric), and $\tau$ is a pre-defined threshold for a prediction to be considered sufficiently close to a ground truth. Note that we define the loss to be 00 when there are no ground truths on an image. Note that the labels $c$ and $\hat{c}$ are also included, and may or not be used (such as to set arbitrarily high values to significantly different classes). We recommend for the sake of consistency to use the same distance as picked for matching, discussed in Section IV-A.
An alternative approach is to the average distance of the predictions from the ground truths.
$L^{\mathrm{cnf}}_{\mathrm{dist}}(\lambda^{\mathrm{cnf}})=\min_{(\hat{b},\hat{c})\in\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)}d\left((b,c),(\hat{b},\hat{c})\right),$
E-B Localization
E-C Classification
E-D Note on Monotonicity
While we mention several times throughout the article the necessity of the loss to be monotone in all its parameters, it is in fact not strictly necessary. Since our optimization algorithm, introduced in the next section, allows to control risks that are non-monotone by monotonizing them on the fly, any losses can be used. However, at the two extremes lie opposite results. On the one hand, when monotonicity is verified, our algorithm offers no penalty. On the other hand, when the loss is strongly non-monotone (think $x\rightarrow x^{2}$), the resulting prediction sets will be overly large and non-informative.
Appendix F Algorithms
In this section, we introduce two algorithms used for optimization of the first- and second-step parameters in our proposed SeqCRC.
Algorithm 3 SeqCRC Monotonizing Optimization for Step 1 (Confidence)
1:Loss functions $L^{\mathrm{cnf}}_{i}$, $L^{\mathrm{loc}}$ and $L^{\mathrm{cls}}$, error level $\alpha^{\mathrm{cnf}}$, Confidence scores $C$
2:Confidence Parameter $\lambda^{\mathrm{cnf}}$
3:Flatten the set of sets of confidence scores $C$ into a single array
4:Create an array of same size, with values the index of image of origin of the associated confidence score in the previous array
5:Sort simultaneously both arrays by decreasing order of confidence scores, denoted by $\mathcal{C}$ and $\mathcal{I}$ respectively.
6:$\lambda^{\mathrm{cnf}}\leftarrow C_{0}$
7:Define a loss array $L$
8:for $i\in 1,\dots,n$ do
9:     Update losses with $\lambda^{\mathrm{cnf}}$ : $L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$, $L^{\mathrm{loc}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$, $L^{\mathrm{cls}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{cls}})$
10:end for
11:Compute risks with updated losses $R^{\mathrm{cnf}}=\sum_{i=1}^{n}L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$, $R^{\mathrm{loc}}=\sum_{i=1}^{n}L^{\mathrm{loc}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$ and $R^{\mathrm{cls}}=\sum_{i=1}^{n}L^{\mathrm{cls}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$,
12:Compute the maximum risk $R=\max\{R^{\mathrm{cnf}},R^{\mathrm{loc}},R^{\mathrm{cls}}\}$
13:for $j\in 2,\dots,\lvert\mathcal{C}\rvert$ do
14:     Update previous risk $R^{\prime}=R$
15:     Update previous $\lambda^{\mathrm{cnf}}{}^{\prime}=\lambda^{\mathrm{cnf}}$
16:     Get current confidence score $c\leftarrow\mathcal{C}(j)$
17:     Get index of the score of the previous image $i\leftarrow\mathcal{I}(j-1)$
18:     Set $\lambda^{\mathrm{cnf}}\leftarrow 1-c$
19:     Update the losses of previous image with new $\lambda^{\mathrm{cnf}}$ : $L^{\mathrm{cnf}}_{i-1}(\lambda^{\mathrm{cnf}})$, $L^{\mathrm{loc}}_{i-1}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$, $L^{\mathrm{cls}}_{i-1}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{cls}})$
20:     Recompute risks with updated losses $R^{\mathrm{cnf}}=\sum_{i=1}^{n}L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$, $R^{\mathrm{loc}}=\sum_{i=1}^{n}L^{\mathrm{loc}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$ and $R^{\mathrm{cls}}=\sum_{i=1}^{n}L^{\mathrm{cls}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$,
21:     Compute the maximum risk $R=\max\{R^{\mathrm{cnf}},R^{\mathrm{loc}},R^{\mathrm{cls}}\}$
22:     if $R<R^{\prime}$ then
23:         $R\leftarrow R^{\prime}$
24:     end if
25:Stopping condition:
26:     if $\frac{n}{n+1}R+\frac{B}{n+1}>\alpha^{\mathrm{cnf}}$ then return previous $\lambda^{\mathrm{cnf}}{}^{\prime}$
27:     end if
28:end for
Algorithm 4 SeqCRC Monotonizing Optimization for Step 2 (Localization/Classification)
1:Loss functions $L^{\bullet}_{i}$, error level $\alpha^{\bullet}$, Confidence scores $C$, a lower $l$ and upper $u$ bound for $\lambda^{\bullet}$, a number of steps $S$, a lower confidence parameter $\lambda^{\mathrm{cnf}}_{-}$
2:Create an empty array $\hat{\Lambda}$
3:for $k\in 1,\cdot,S$ do
4:     $\lambda^{\bullet}\leftarrow\frac{l+u}{2}$
5:     Flatten the set of sets of confidence scores $C$ into a single array
6:     Create an array of same size, with values the index of image of origin of the associated confidence score in the previous array
7:     Sort simultaneously both arrays by decreasing order of confidence scores, denoted by $\mathcal{C}$ and $\mathcal{I}$ respectively.
8:     $\lambda^{\mathrm{cnf}}\leftarrow C_{0}$
9:     Define a loss array $L$
10:     for $i\in 1,\dots,n$ do
11:         Update $\Lambda^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$
12:     end for
13:     Compute risk $R=\sum_{i=1}^{n}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}})$
14:     for $j\in 2,\dots,\lvert\mathcal{C}\rvert$ do
15:         Update previous risk $R^{\prime}=R$
16:         Update previous $\lambda^{\mathrm{cnf}}{}^{\prime}=\lambda^{\mathrm{cnf}}$
17:         Get current confidence score $c\leftarrow\mathcal{C}(j)$
18:         Get index of the score of the previous image $i\leftarrow\mathcal{I}(j-1)$
19:         Set $\lambda^{\mathrm{cnf}}\leftarrow 1-c$
20:         Update the loss of previous image with new $\lambda^{\mathrm{cnf}}$ : $L^{\bullet}_{i-1}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$
21:         Recompute risk with updated loss $R=\sum_{i=1}^{n}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$
22:         if $R<R^{\prime}$ then
23:              $R\leftarrow R^{\prime}$
24:         end ifStopping condition:
25:         if $\lambda^{\mathrm{cnf}}=\lambda^{\mathrm{cnf}}_{-}$ then
26:              break
27:         end if
28:     end for
29:     if $\frac{n}{n+1}R+\frac{B}{n+1}\leq\alpha^{\bullet}$ then
30:         Append $\lambda^{\bullet}$ to $\hat{\Lambda}$
31:         $u\leftarrow\lambda^{\bullet}$
32:     else
33:         $l\leftarrow\lambda^{\bullet}$
34:     end ifreturn The last element of $\hat{\Lambda}$
35:end for
TABLE V: Experimental Results using the DETR-50 model for $\alpha_{tot}=0.1$

### v2

@@ A-B A simple symmetry property
Appendix B Additional Losses
As discussed in details in Appendix C, many natural image-level losses are aggregations of object-level losses. While there are exceptions to that, the losses presented in this appendix are of this form. Formally, denote by $\Phi:S(\mathbb{R})\mapsto\mathbb{R}$ an aggregation function, where $S(\mathbb{R})$ correspond to the set of all the finite sequence of real values. We introduce three main aggregation strategies.
  1. 1.
  2. 2.
    Max: $\Phi_{\mathrm{max}}(S)=\max_{i}S_{i}$
  3. 3.
    Thresholded Average (at level $\tau$): $\Phi_{\mathrm{thr}}(S)=\mathbbm{1}_{\Phi_{\mathrm{avg}}(S)>\tau}$
B-A Confidence
where $d:(\mathcal{B}\times\mathcal{C})\times(\mathcal{B}\times\Sigma_{K-1})\rightarrow\mathbb{R}^{+}$ is any function measuring a distance (particularly, it does not need to be symmetric), and $\tau$ is a pre-defined threshold for a prediction to be considered sufficiently close to a ground truth. Note that we define the loss to be 0 when there are no ground truths on an image. Note that the labels $c$ and $\hat{c}$ are also included, and may or not be used (such as to set arbitrarily high values to significantly different classes). We recommend for the sake of consistency to use the same distance as picked for matching, discussed in Section IV-A.
Alternatively, we can average the distance between the predictions and the ground truths. This is a non-binarized version of the loss function mentioned above.
[…]
B-B Localization
B-C Classification
B-D Note on Monotonicity
While we mention several times throughout the article the necessity of the loss to be monotone in all its parameters. As mentioned in Remark 1, and introduced in  IV-E, we apply a monotonization process that alter the considered losses to make them compatible with our theoretical results. Since our optimization algorithm, introduced in the next section, allows to control losses that are non-monotone by monotonizing them on the fly, any losses can be used. However, at the two extremes lie opposite results. On the one hand, when monotonicity is verified, our algorithm offers no penalty. On the other hand, when the loss is strongly non-monotone
, the resulting prediction sets will be overly large and non-informative.
Appendix C Types of guarantee & Matching
Types of guarantee:
On a given image, there can a priori be any number of objects (and their associated bounding boxes). On the one hand, each of those individual objects may be considered as ground truth individually, i.e. $Y_{i}=(b_{i},c_{i})$. On the other hand, the ground truth can be defined as the set of all objects present on an image, i.e. $Y_{i}=(b_{i,j},c_{i,j})_{j=1}^{n_{i}}$, as introduced in this work. Respectively, we refer to those as object-level and image-level guarantees following [15] as well as  [14, 16] where object-level guarantees have previously been referred to as box-level. However, the exact setting of the latter varied, as in these previous works, conformal prediction was only applied at the object level to true-positive bounding boxes (i.e., ground-truth - prediction pairs, with only pairs with IoU above a certain threshold $\beta$ considered).
The choice between object-level and image-level setting depends on practical requirements and should be reflected in the definition of the losses. An example is a binary loss: in the object-level setting, it entails that a proportion at least $1-\alpha$ of ground truth bounding boxes are correctly covered, while in the image-level setting, it entails that, on a proportion $1-\alpha$ of images at least, all bounding boxes (or a desired proportion) are covered.
Matching:
As common in object detection [40, 43], a matching between the true bounding boxes and predicted ones is necessary, not only for evaluation here but also for calibration. We refer to by matching, a family of functions $\pi_{i}$ defined for each image $X_{i}$ that maps the index $j$ of an object on $X_{i}$ to the index of a predicted box among $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})$. More formally, $\pi_{i}:\{1,\cdots,n_{i}\}\rightarrow\{1,\cdots,\lvert\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})\rvert\}$ with $n_{i}$ the number of objects on image $X_{i}$ and $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})$ the sets of bounding boxes likely to correspond to objects on image $X_{i}$. We do not make further assumption on the matching algorithm. Therefore it includes injective functions, i.e. each prediction being matched to at most one ground truth. We can also consider functions where some inputs might have identic output, i.e. reusing predictions for multiple ground truth. An example of an injective matching is the commonly used Hungarian algorithm.
Appendix D Algorithms
In this section, we introduce two algorithms to compute the first- and second-step parameters of our SeqCRC method.
Recall that we use a monotonization trick to enforce monotonicity of $L^{\mathrm{loc}}$ and $L^{\mathrm{cls}}$ in $\lambda^{\mathrm{cnf}}$. That is, we replace $L^{\mathrm{loc}}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}})$ and $L^{\mathrm{cls}}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}})$ with $\sup_{\lambda^{\prime}\geq\lambda^{\mathrm{cnf}}}L^{\mathrm{loc}}_{i}(\lambda^{\prime},\lambda^{\mathrm{loc}})$ and $\sup_{\lambda^{\prime}\geq\lambda^{\mathrm{cnf}}}L^{\mathrm{cls}}_{i}(\lambda^{\prime},\lambda^{\mathrm{cls}})$ respectively. For computational and memory efficiency purposes, these suprema are computed on the fly (i.e., while optimizing the underlying risks).
More precisely, the pseudo-code of Algorithm 3 can be used to compute $\lambda^{\mathrm{cnf}}_{+}$ and $\lambda^{\mathrm{cnf}}_{-}$ defined in (4) and (5), where $R^{\mathrm{loc}}_{n}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$ and $R^{\mathrm{cls}}_{n}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{cls}})$ (the localization and classification components of $\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})$, see (3)) are obtained after monotonizing the localization and classification losses with respect to $\lambda^{\mathrm{cnf}}$.For broader applicability, in the actual code, we also monotonize the losses $L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$ within $R^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})$ in (3), even if all $L^{\mathrm{cnf}}$ examples given in Section IV-B are already monotone. To implement the monotonization trick, we use the key fact that the set $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$ defined in (11) is a piecewise-constant function of $\lambda^{\mathrm{cnf}}$ (in a right-continuous manner), with discontinuity points given by the confidence scores $o(x)_{[k]}$ of predicted objects.
Similarly, the pseudo-code of Algorithm 4 shows how to compute $\lambda^{\mathrm{loc}}_{+}$ and $\lambda^{\mathrm{cls}}_{+}$ defined in (6), where the risk $R^{\bullet}_{n}(\lambda^{\mathrm{cnf}}_{-},\lambda^{\bullet})$ is obtained after monotonizing the localization or classification losses with respect to $\lambda^{\mathrm{cnf}}$. More precisely, we output (slight) upper-approximations of $\lambda^{\mathrm{loc}}_{+}$ and $\lambda^{\mathrm{cls}}_{+}$ using a binary search.
Both algorithms are written for $\Lambda^{\mathrm{cnf}}=[0,1]$.
Algorithm 3 SeqCRC optimization for Step 1 (confidence), with monotonization trick
1:Loss functions $(L^{\mathrm{cnf}}_{i})_{i=1}^{n}$, $(L^{\mathrm{loc}}_{i})_{i=1}^{n}$ and $(L^{\mathrm{cls}}_{i})_{i=1}^{n}$, error level $\alpha^{\mathrm{cnf}}$, loss bound $B$, parameter bounds $\bar{\lambda}^{\mathrm{loc}},\bar{\lambda}^{\mathrm{cls}}$, confidence scores $C$ over whole dataset (list of $n$ sublists).
2:Confidence parameter $\lambda^{\mathrm{cnf}}_{+}$ (if input $B$ equals $\tilde{B}^{\mathrm{cnf}}$) or $\lambda^{\mathrm{cnf}}_{-}$ (if $B=0$); see (4) and (5).
3:Flatten the hierarchical list $C$ of confidence scores into a single array. Create an array II of the same size, filled with the associated image indices.
4:Sort both arrays simultaneously by increasing order of confidence scores, and denote the results by $\mathcal{C}$ and $\mathcal{I}$.
5:Shift $\mathcal{C}$’s values to the left, and replace the last (missing) element by $1$.
6:% Next we search for the infimum $\lambda^{\mathrm{cnf}}$ value in (4) or (5), starting from $\lambda^{\mathrm{cnf}}=1$ downwards, while computing the losses $L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$ and the monotonized losses $\sup_{\lambda^{\prime}\geq\lambda^{\mathrm{cnf}}}L^{\bullet}_{i}(\lambda^{\prime},\bar{\lambda}^{\bullet})$ on the fly.
[…]
8:Define three (monotonized) loss arrays: $\mathrm{Lcnf}$, $\mathrm{Lloc}$, $\mathrm{Lcls}$.
9:for $i=1,\dots,n$ do
10:  Compute losses with $\lambda^{\mathrm{cnf}}$: $\mathrm{Lcnf}[i]\leftarrow L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$, $\mathrm{Lloc}[i]\leftarrow L^{\mathrm{loc}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$, and $\mathrm{Lcls}[i]\leftarrow L^{\mathrm{cls}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{cls}})$.
11:end for
12:Compute the three risks $R^{\mathrm{cnf}}=n^{-1}\sum_{i=1}^{n}\mathrm{Lcnf}[i]$, $R^{\mathrm{loc}}=n^{-1}\sum_{i=1}^{n}\mathrm{Lloc}[i]$ and $R^{\mathrm{cls}}=n^{-1}\sum_{i=1}^{n}\mathrm{Lcls}[i]$.
13:Compute the maximum risk $R=\max\{R^{\mathrm{cnf}},R^{\mathrm{loc}},R^{\mathrm{cls}}\}$.
14:% We now iteratively decrease the value of $\lambda^{\mathrm{cnf}}$.
15:for $c,i\in\mathcal{C},\mathcal{I}$ do
16:  Store previous value $\lambda^{\mathrm{cnf}}{}^{(\mathrm{old})}\leftarrow\lambda^{\mathrm{cnf}}$.
17:  Set $\lambda^{\mathrm{cnf}}\leftarrow 1-c$.
18:  On image $i$, update confidence loss and monotonized localization/classification losses:
[…]
22:  Compute the three risks $R^{\mathrm{cnf}}=n^{-1}\sum_{i=1}^{n}\mathrm{Lcnf}[i]$, $R^{\mathrm{loc}}=n^{-1}\sum_{i=1}^{n}\mathrm{Lloc}[i]$ and $R^{\mathrm{cls}}=n^{-1}\sum_{i=1}^{n}\mathrm{Lcls}[i]$.
23:  Compute maximum risk $R=\max\{R^{\mathrm{cnf}},R^{\mathrm{loc}},R^{\mathrm{cls}}\}$.
24:  % Stopping condition:
25:  if $\frac{n}{n+1}R+\frac{B}{n+1}>\alpha^{\mathrm{cnf}}$ then return previous $\lambda^{\mathrm{cnf}}{}^{(\mathrm{old})}$.
26:  end if
27:end for
28:return 0 % if stopping condition is never satisfied
Algorithm 4 SeqCRC optimization for Step 2 (localization/classification), with monotonization trick
1:Loss functions $(L^{\bullet}_{i})_{i=1}^{n}$, error level $\alpha^{\bullet}$, loss bound $B^{\bullet}$, confidence parameter $\lambda^{\mathrm{cnf}}_{-}$ from Step 1, confidence scores $C$ over the whole dataset (list of $n$ sublists), a priori lower and upper bounds $l,u$ on $\lambda^{\bullet}_{+}$, number $S$ of steps.
2:Localization/classification parameter $\lambda^{\bullet}_{+}$; see (6).
3:Create an empty array $\hat{\Lambda}$
4:for $k=1,\ldots,S$ do
5:  $\lambda^{\bullet}\leftarrow\frac{l+u}{2}$
6:  Flatten the hierarchical list $C$ of confidence scores into a single array. Create an array II of the same size, filled with the associated image indices.
7:  Sort both arrays simultaneously by increasing order of confidence scores, and denote the results by $\mathcal{C}$ and $\mathcal{I}$.
8:  Shift $\mathcal{C}$’s values to the left, and replace the last (missing) element by $1$.
9:  % Next we compute the monotonized losses $\sup_{\lambda^{\prime}\geq\lambda^{\mathrm{cnf}}_{-}}L^{\bullet}_{i}(\lambda^{\prime},\lambda^{\bullet})$, starting from $\lambda^{\prime}=1$ downwards.
10:  Define a monotonized loss array $L$.
11:  $\lambda^{\prime}\leftarrow 1$
12:  for $i=1,\dots,n$ do
13:   Set $L[i]\leftarrow L^{\bullet}_{i}(\lambda^{\prime},\lambda^{\bullet})$
14:  end for
15:  for $c,i\in\mathcal{C},\mathcal{I}$ do
16:   Set $\lambda^{\prime}\leftarrow 1-c$
17:   Update monotonized loss of image $i$ using confidence parameter $\lambda^{\prime}$: $L[i]\leftarrow\max\{L[i],L^{\bullet}_{i}(\lambda^{\prime},\lambda^{\bullet})\}$
18:   if $\lambda^{\prime}\leq\lambda^{\mathrm{cnf}}_{-}$ then % Stopping condition
19:     break
20:   end if
21:  end for
22:  Compute the risk with monotonized losses: $R\leftarrow n^{-1}\sum_{i=1}^{n}L[i]$
23:  if $\frac{n}{n+1}R+\frac{B^{\bullet}}{n+1}\leq\alpha^{\bullet}$ then
24:   Append $\lambda^{\bullet}$ to $\hat{\Lambda}$
25:   $u\leftarrow\lambda^{\bullet}$
26:  else
27:   $l\leftarrow\lambda^{\bullet}$
28:  end if
29:return the last element of $\hat{\Lambda}$ if nonempty; otherwise, raise an error
30:end for
Appendix E Details on Experiments
@@
E-A Complete benchmark
We include the complete results of our experiments here in Tables  VII, VIII, V, and VI. All experiments are conducted on a single NVIDIA RTX 4090. Each individual experiment takes approximately 20 minutes of runtime, most of it being taken by the monotonizing optimization algorithm. We now introduce multiple additional visualizations of our best-performing experiment run, for additional discussions on the limitations of conformal object detection. On Fig. 9, we observe two successful detections, with singleton classification prediction sets and small localization margins. However, on the third image, the moon is predicted to be a kite. While this is incorrect due to the monotonicity constraints of CRC, we may only provide recall guarantees, i.e., ensure that we do not miss the ground truth. In that framework, additional predictions such as the moon predicted as a kite, are not penalized. In Fig. 11, we observe another instance of the previous mistake, since the boat is detected twice: once excluding its mast, and once including it. The next two images illustrate two failure cases of our guarantee, corresponding to previously discussed issues in terms of abuses of labeling. On the middle image, books are sometimes annotated individually, and sometimes as a block. On the right image, a statue is annotated as a person. The variability in the definition of ground truths leads to perceived failures, that are not necessarily actual failures. As visible on both image (b) and (c), the clock (bottom-right) and statue (person, top) are respectively assigned a book and a clock predictions by our matching. In order to reach lower levels of error rates $\alpha$, these mistakes would have to be controlled, and would result in a need for significant margins at inference (or models that are able to learn this bias).
E-B Additional Experiments
@@
E-B1 mAP vs CP
@@ E-B2 Margins
N. margins Value of the Margin Cov Mean Set Size
@@
Figure 12: Histogram of the number of objects per image for Ground Truths (orange), Predictions with only pre-filtering (Blue), and Prediction after confidence threshold (with the Box Count Recall loss, $\alpha^{\mathrm{cnf}}=0.02$ and DETR-50)
E-B3 Distribution of Object Counts
In this section we visualize on Fig. 12 the empirical distribution of the number of objects for the ground truths and predictions. The predictions are obtained using DETR-101 and are displayed in two variations. The blue histogram corresponds to the number of objects only after filtering out those below $0.001$ confidence, and the green histogram using a confidence threshold obtained for the box count recall loss with $\alpha^{\mathrm{cnf}}=0.02$. First, notice that the number of predicted objects can be quite large, and is heavy-tailed. We then notice that while the obtained confidence threshold $0.074$ appears quite low, the distribution of counts seems to get closer to the ground truth distribution (plotted in orange). Indeed, it remains conservative with regards to the ground truth object counts generally, while not being too loose in many cases. Additionally, while choosing the standard threshold $0.5$ leads to a distribution (not displayed here for clarity) matching the ground truth one even more, it is not guaranteed to contain, with high probability, enough objects.
Appendix F Effect of Object Detection and Conformal Settings
@@
Lastly, regarding the $x_{1}y_{1}x_{2}y_{2}$ (coordinates of two corners) versus xywhxywh (coordinates of the center + width and height) bounding box representations, while fundamentally equivalent, their application within CP algorithms yields divergent results. The xywhxywh format inherently involves symmetric corrections and a translational adjustment, whereas the xyxyxyxy format necessitates independent corrections for each side. However, these approaches can be reformulated to yield equivalent predictive sets.
TABLE V: Experimental Results using the DETR-101 model for $\alpha^{\mathrm{tot}}=0.1$